OpenStudy (liizzyliizz):

for f(x)= csc piex ( there would be more then one vertical asymptote right?) now how would u write that...)

OpenStudy (anonymous):

\[\frac{1}{\sin(\pi x)}\] asymptote where \[\sin(\pi x)=0\] so any integer would make it zero

OpenStudy (anonymous):

since \[\sin(k\pi)=0\] for any integer k